### v1

# Constrained Black-Box Attacks Against Multi-Agent Reinforcement Learning

Abstract
Collaborative multi-agent reinforcement learning (c-MARL) has rapidly evolved, offering state-of-the-art algorithms for real-world applications, including sensitive domains. However, a key challenge to its widespread adoption is the lack of a thorough investigation into its vulnerabilities to adversarial attacks. Existing work predominantly focuses on training-time attacks or unrealistic scenarios, such as access to policy weights or the ability to train surrogate policies. In this paper, we investigate new vulnerabilities under more realistic and constrained conditions, assuming an adversary can only collect and perturb the observations of deployed agents. We also consider scenarios where the adversary has no access at all. We propose simple yet highly effective algorithms for generating adversarial perturbations designed to misalign how victim agents perceive their environment. Our approach is empirically validated on three benchmarks and 22 environments, demonstrating its effectiveness across diverse algorithms and environments. Furthermore, we show that our algorithm is sample-efficient, requiring only 1,000 samples compared to the millions needed by previous methods.
Index Terms:
Multi-agent reinforcement learning, adversarial attack, black box attack
I Introduction
Collaborative multi-agent reinforcement learning (c-MARL) algorithms have demonstrated state-of-the-art performances in complex cooperative tasks [26, 29], making them well-suited for solving real-world problems across various domains [14, 21, 32]. However, a critical prerequisite for the widespread adoption of c-MARL is a full understanding of its vulnerabilities to adversarial attacks [7, 9], particularly when deployed.
While much of the literature on adversarial attacks against c-MARL focuses on training-time attacks [33, 13, 2, 6], we focus instead on test-time attacks, where the adversary is present during deployment. Prior work on test-time attacks [23, 11, 17] has primarily considered white-box threat models, in which the adversary has access to the policy architecture and its parameters. This scenario is not always feasible. Moreover, we argue that having such access can be considered a successful attack in itself, as it typically represents proprietary knowledge with significant financial implications if leaked. In contrast, black-box threat models [7] do not assume access to the policy’s weights or its architecture; instead, they often involve learning a surrogate policy network. This can be achieved either by training the policy from scratch [7] or through imitation learning [28, 8]. The former requires access to the training environment, while the latter relies on having access to both observations and actions (i.e., expert demonstrations) or the ability to query the model. But what if the adversary is prevented from having such access to train the surrogate policy?
In our work, we examine test-time attacks against c-MARL in black-box settings, but we push the standard assumptions a step further: we consider an adversary whose access is restricted to the observations (no access to the actions) of deployed agents, referred to as infected agents. The adversary does not have full control over these infected agents; instead, they can only add small perturbations to their observations. To the best of our knowledge, we are the first to investigate adversarial attacks under such constrained conditions.
Figure 1: Assumption Strength vs Plausibility: There is an inverse relationship between the strength of a threat model’s assumptions and its practical relevance: the less access an attacker is assumed to have, the more plausible the scenario becomes in real-world settings.
The first main contribution of our paper is to answer the question of whether an adversary with such limited information and capabilities can sabotage a c-MARL system and, if so, how they might achieve this. Not only is the answer yes, but the adversary can inflict significant damage with as few as 1,000 collected samples, in stark contrast to previous approaches requiring millions of samples [23].
We propose a novel, simple, and yet very effective algorithm that will stealthily and strategically manipulate the observations of compromised agents. The goal is to add small perturbations to these observations, causing the agents to perceive the same environment differently- a phenomenon we refer to as misalignment. For instance, imagine a simplified version of the Pursuit game [5], where two agents must cooperate to surround a moving object to receive a reward. If each agent receives conflicting information about the object’s position due to an adversarial attack, their ability to coordinate becomes severely compromised. Our crafted perturbations are designed specifically to induce such misalignment to the targeted agents. We refer to this attack as Align attack.
The second main contribution of our paper is to propose an even more constrained adversarial attack: we consider an adversary with no access whatsoever (no observations, actions, or policies), possessing only the capability to inject small perturbations into each agent’s observation stream. The literature typically refers to this type of attack as Free attacks, which predominantly consist of injecting random noise. However, in our paper, we use the same idea of introducing misalignment in agents and make use of partial Hadamard matrices to generate better perturbations. As we explain later, misalignment can be induced using orthogonal matrices. We refer to this attack as Hadamard attack.
Additionally, we propose a targeted attack that combines the two approaches: we use observations to select a subset of critical agents, and then use Hadamard matrices to craft adversarial perturbations. This allows us to benefit from the profiling capability of the Align attack and the fast generation of the Hadamard attack, thereby resulting in an efficient and lightweight attack.
Our contributions are illustrated in Figure 1. Most previous work considered attacks that rely on having access to some form of policy: the deployed policy (i.e., white-box attacks), a surrogate policy trained from scratch in the same environment, or a policy trained using imitation learning. These approaches depend on the transferability [19] of adversarial examples and tend to be very effective. However, as previously mentioned, they rely on strong assumptions about the attacker’s access, which reduces their plausibility. In contrast, we focus on threat models that assume weak access, which we believe are more realistic and yet remain underexplored in the literature. We argue that, in general, the fewer elements the attacker is assumed to access, the more plausible the attack becomes. Having access to observations only is more plausible than having access to both observations and actions, and both are more realistic than assuming access to the policy. Assuming no access at all is the most plausible scenario. We can summarize this intuition with the following relationship (The curve shown in Figure 1):
$\text{Plausibility}\approx\frac{1}{\text{Assumption strength}}$ (1)
It is important to note, however, that assuming no access does not mean the attack is trivial to implement. Like all the other threat models, we still assume the attacker is somehow present in the system and can inject perturbations. We only focus on what information the attacker relies on to generate those perturbations.
Finally, we demonstrate the effectiveness of our approach across three MARL benchmarks and 22 different tasks, covering fully-observable, partially-observable, and highly cooperative tasks. Our results show that our adversaries significantly undermine the performance of collaborative agents. Moreover, our attack applies to both value-based and policy-based algorithms. We also present a detailed analysis of the key factors that are particularly important for the performance and the computational cost of our attacks.
II Related works
Adversarial reinforcement learning can be classified according to several criteria: Is the system composed of a single agent or multiple agents? Does the attack occur during training or deployment? Are we considering a white-box or black-box scenario? What system components does the attacker have access to, and which does it target? Most existing work focuses on adversarial attacks against single-agent systems [18, 7, 8, 12, 22, 9, 28], often under white-box settings with access to states, actions, and rewards. In contrast, our work addresses adversarial attacks on deployed multi-agent systems in a black-box setting. We consider scenarios where the attacker initially has access only to the agents’ observations, and further extend our approach to cases where the adversary has no access at all. Table I provides a comprehensive comparison between our work and previous work on adversarial c-MARL.
TABLE I: Comparison of our work and previous papers. We indicate whether the work considers test-time vs training-time attacks and black-box vs white-box settings. We report the attacker’s accessible elements and targeted components.
Paper Test-time Black-box Accessible Elements Target
Policy Obs Actions Reward Env Actions Obs Reward Env
Pham et al. (2023) [23]
Lin et al. (2020) [11]
Nisioti et al. (2021) [17]
Hu and Zhang (2022) [6]
Chen et al. (2024) [2]
Liu and Lai (2023) [13]
Zheng et al. (2023) [33]
Ours (✓,✗)
Training-time vs Test-time attacks. Training-time attacks, also referred to as data poisoning attacks [24], occur when an adversary is present during the training. The aim of these attacks is to manipulate the agent into learning a target policy crafted by the attacker. This can be accomplished through reward poisoning [13, 31] or environment poisoning [25], as well as by manipulating observations or actions [6, 2, 33]. However, it is not always feasible for the adversary to interfere with the training process. Multi-agent systems are arguably more vulnerable when deployed in the real world. Attacks occurring during deployment are referred to as test-time attacks, where the goal is to degrade the performance of already-trained policies. This is typically achieved by exploiting the known vulnerabilities of neural networks to adversarial inputs [7, 27], mainly through observation manipulation [23, 11], but also through action manipulation [17]. Our work focuses on test-time attacks that target observations. Much of prior work (see Table I) on test-time attacks assumes access to numerous elements at the same time, such as the policy network, observation, actions, rewards, and even the training environment—assumptions that are unrealistic in actual deployment scenarios. Therefore, we extend previous work by considering more practical scenarios, where the adversary has limited access, either only to the agents’ observations or no access at all.
White-box vs Black-box. In white-box scenarios, the attacker knows the learning algorithm and has access to the policy weights and its architecture [2, 6, 17, 11, 23, 7]. While these attacks tend to be the most effective, it is impractical for the attacker to have the complete knowledge of the deployed policies. Conversely, black box settings allow for more relaxed assumptions [7]. Most existing work relies on learning a surrogate policy to exploit the transferability of adversarial examples [19]. This surrogate policy can be learned by training the model from scratch [7], but doing so requires access to the training environment. Alternatively, imitation learning can be used to approximate the policy [28, 8], which would necessitate access to observation-action pairs or the ability to query the deployed policy. However, neither of these approaches is applicable in our scenario, as we lack access to the training environment and actions, and we cannot query the model.
Finally, as shown in Table I, we are the only work that investigates test-time attacks in black-box settings with minimal accessible information.
III Background
Consider a neural network $f$ parameterized by $\theta$, which takes an input $x\in\mathcal{X}$ and outputs $y$. The network is trained to minimize a loss function J. The goal of an adversarial attack is to add a perturbation $\delta$ to the original input $x$ in order to maximize the loss $\text{J}(x+\delta,y;\theta)$ with respect to $\delta$. Moreover, this perturbation must be small enough to avoid detection, which can be achieved by constraining the $L_{\infty}$ norm of the perturbation within an attack budget $\epsilon$. The perturbation $\delta$ can be determined by solving the following optimization problem:
$\delta=\arg\max_{\delta}\text{J}(x+\delta,y;\theta)\quad\text{subject to}\quad\|\delta\|_{\infty}\leq\epsilon$ (2)
A wide range of attack methods has been proposed for generating adversarial perturbations. Fast Gradient Signed Methods (FGSM) [4] is a straightforward technique that generates perturbations as follows:
$\delta=\epsilon\times\text{sign}(\nabla_{x}(\text{J}(x,y;\theta)))$ (3)
where $\nabla_{x}(\text{J}(x,y;\theta))$ is the gradient of the loss function J with respect to the input $x$ and sign is a function which returns +1 if the argument is positive, -1 if negative, and 0 if zero.
FGSM is a single-step attack. A more powerful variant is the multistep attack known as Projected Gradient Descent (PGD) [10, 15]. This attack iteratively injects perturbations over $K$ steps with a small step size $\alpha$. Moreover, to ensure that the perturbed input remains within the valid input domain $\mathcal{X}$, the perturbed data is projected back into the domain after each step. PGD perturbs the inputs as follows:
$x^{*}_{0}=x,x^{*}_{t+1}=Clip_{\mathcal{X}}\{x^{*}_{t}+\alpha\times\text{sign}(\nabla_{x}\text{J}(x^{*}_{t},y;\theta))\}$ (4)
where $Clip_{\mathcal{X}}$ is an element-wise clipping operator that ensures the results remain within the valid domain $\mathcal{X}$.
IV Methodology
IV-A Problem statement
We consider a cooperative multi-agent system during its deployment phase. Let $\mathcal{N}=\{1,\dots,n\}$ be the set of agents. We assume that an adversary has successfully compromised the entire set $\mathcal{N}$ or a subset of it. Although the adversary has compromised these agents, it does not possess direct control over their actions. Instead, for each compromised agent $i\in\mathcal{N}$, the adversary can inject adversarial perturbations before they are processed by the agent’s policy. The attacker must ensure that these perturbations remain small enough to avoid detection.
The main problem we seek to address can be outlined as follows:
Given the local observations of the agents, can an adversary undermine the performance of a deployed cooperative multi-agent system? And if these observations are inaccessible, can we still design effective noise-based adversarial attacks?
In the following paragraphs, we first explain the intuition behind our paper and then present our adversarial algorithms.
IV-B Our intuition
In cooperative tasks, the ability of agents to work together relies heavily on having aligned perceptions and beliefs of their environment. By aligned perceptions, we mean that agents observing the same object must perceive it consistently, which includes receiving identical information about its characteristics. For example, in the SMAC games [3], agents work together to defeat a set of enemy units in different battles. To succeed, agents must learn a range of important skills such as focus fire, where agents coordinate to jointly attack a single opponent to eliminate it quickly (see Figure 2). This requires the agents to agree on which enemy to jointly attack, typically selecting the weakest one (i.g, based on the health level). Thus, to properly use this skill, the team members should all have aligned perceptions of their targeted opponent, including its health and position. If this alignment is lacking, due to the adversary tampering with health values as in Figure 2, agents may fail to coordinate their attacks. This illustrates how misaligned perceptions can hinder agents’ ability to collaborate effectively and achieve shared goals, potentially pushing them to take opposing actions. Thus, if an adversary manages to properly manipulate agents’ perceptions, the collaboration is likely to be compromised.
We do not necessarily assume that the agents must observe the same objects. The same intuition applies to partially observable environments where there is minimal to no overlap between agents’ observations. In these scenarios, our intuition is to induce misalignment in the common beliefs among the agents [30, 16]. Accordingly, our experimental section mainly focuses on partially observable tasks.
To summarize, the intuition behind our attack is to exploit this reliance on aligned perceptions and understandings among agents. We aim to perturb the observations of infected agents in a way that induces misalignment in their perceptions. This will create confusion and disagreements in how they understand their environment. To achieve this, we adopt a divide-and-conquer strategy by ensuring that each infected agent perceives the environment differently from the other compromised agents. The goal is to divide the compromised agents by creating discrepancies in their perceptions, severely undermining their ability to coordinate, which ultimately weakens the overall performance of the entire team.
In the following sections, we introduce two novel attacks. Sections IV-C and IV-D present the Align attack and its lightweight variant, respectively, while Section IV-E describes the Hadamard attack.
IV-C The Align attack
Figure 2: Illustration: a small adversarial perturbation (0.01,0,0.04)(0.01,0,-0.04) could prevent the agents from coordinating their actions effectively
The standard assumption in c-MARL systems is that collaborative agents typically have aligned perceptions, meaning they usually receive consistent information from their environments. Therefore, it is accepted to assume that the observations of the infected agents $\{o_{1},o_{2},\dots,o_{n}\}$ are aligned. Our goal is to manipulate each observation so that the new observations $\{o_{1}+\delta_{1},o_{2}+\delta_{2},\dots,o_{n}+\delta_{n}\}$ become misaligned, resulting in conflicting information among those agents. However, before we can do this, we must first be able to measure whether the observations are misaligned, as this will inform our attack.
When agents have aligned perceptions, their observations tend to correlate because they contain similar information or have comparable understandings of the environment. This means that the observations of one agent can be approximated based on the observations of other agents. On the other hand, when their observations are misaligned, they become less correlated and contain less overlapping content.
Formally, when agents’ perceptions are aligned, we can train the following neural network $f_{\theta}$:
$f_{\theta}(o_{-i})\approx o_{i}\quad\forall i\in\mathcal{N}$ (5)
where $o_{-i}=\{o_{1},\dots,o_{i-1},o_{i+1},\dots o_{n}\}$ is the concatenated observations of all the agents except agent $i$. This network can be trained using the mean squared error:
$\text{J}(\mathbf{o};\theta)=\frac{1}{n}\sum_{i\in\mathcal{N}}\left(f_{\theta}(o_{-i})-o_{i})\right)^{2}$ (6)
where $\mathbf{o}$ is the joint observation.
Once $f_{\theta}$ is properly trained, it can effectively reconstruct each agent’s observations using the observations from its peers, which will manifest as a low loss value. Conversely, when the network is fed with misaligned observations, we will notice high loss values. Therefore, a simple strategy for the adversary to induce misalignment to the agents is to add a small perturbation $\{\delta_{i}\}_{i\in\mathcal{N}}$ to the observations that will maximize the loss.
This strategy can be formally expressed as:
$\delta=\arg\max_{\delta}\text{J}(\mathbf{o}+\delta;\theta)\quad\text{s.t}\quad\|\delta\|_{\infty}\leq\epsilon$ (7)
Our problem in Equation (7) is similar to the one in Equation (2), suggesting that we can use PGD. However, there is a subtle detail that must be considered: in Equation (2) we add the perturbation to the input, whereas in Equation (7) we add it to both the input and the output.
In light of these factors, our attack can be split into the following two phases, as summarized in Algorithm 1:
Phase One: data collection. The purpose of this step is to collect a dataset $\mathcal{D}$ to train the neural network $f_{\theta}$. This dataset is gathered by storing the observations of infected agents during test time. Once we have accumulated observations for a sufficient period $\mathcal{T}^{c}$, we train the neural network $f_{\theta}$. This network is trained only once, and its parameters are frozen after this phase. No attack is executed during this phase.
Phase Two: adversarial attack. In this phase, we launch our adversarial attack. At each time step $t$, we start by intercepting the current observations $\{o_{i,t}\}_{1,\dots,n}$. Next, we generate the corresponding perturbations using PGD. Our exact implementation of PGD is outlined in Algorithm 2. Following this, we inject these perturbations into the respective agents.
Algorithm 1 Adversarial Attack to Induce Misalignment
1:Input: Data collection period $\mathcal{T}^{c}$, attack period $\mathcal{T}^{a}$, PGD parameters.
2:Initialization: Initialize $f_{\theta}$ and the dataset $\mathcal{D}=\varnothing$
3:Phase One: Data collection and model training
4:for $t=1$ to $\mathcal{T}^{c}$  do
5:  Collect $\{o_{1,t},o_{2,t},\dots,o_{n,t}\}$ and store it in $\mathcal{D}$.
6:end for
7: Train $f_{\theta}$ using the collected dataset $\mathcal{D}$
8:Phase Two: Adversarial attack on infected agents
9:for  $t=1$ to $\mathcal{T}^{a}$ do
10:  Compute the perturbation $\delta$ using PGD (Algorithm 2)
11:  Inject the perturbations
12:end for
Algorithm 2 Projected Gradient Descent algorithm
1:Input: Trained neural network $f$, observations $\mathbf{o}$, attack budget $\epsilon$, step size $\alpha$, number of iteration $K$, loss function J, allowed observation values $o^{min},o^{max}]$.
2:Initialization: Perturbation $\delta^{(0)}=0$, perturbed observation $\tilde{\mathbf{o}}^{(0)}=\mathbf{o}$.
3:Output: Perturbed observations: $\tilde{\mathbf{o}}^{(0)}=\tilde{\mathbf{o}}^{(K)}$ .
4:for  $k=1$ to $K$ do
5:  Compute the gradient $\nabla_{\tilde{\mathbf{o}}^{(k-1)}}\text{J}(\tilde{\mathbf{o}}^{(k-1)};\theta)$
6:  Compute $\delta^{k}$: $\delta^{(k)}=\alpha\times\text{sign}(\nabla_{\tilde{\mathbf{o}}^{(k-1)}}\text{J}(\tilde{\mathbf{o}}^{(k-1)};\theta))$
7:  Clip $\delta^{(k)}$ to $\epsilon,\epsilon]$
8:  Compute $\tilde{\mathbf{o}}^{(k)}$ $\tilde{\mathbf{o}}^{(k)}=\tilde{\mathbf{o}}^{(k-1)}+\delta^{(k)}$
9:  Clip $\tilde{\mathbf{o}}^{(k)}$ to $o^{min},o^{max}]$
10:end for
IV-D A lighter Align attack
It is not always practical to compute and inject adversarial perturbations for each agent. In addition to avoiding detection, one of the main reasons is the associated computational cost. A straightforward approach for handling this is to limit the number of agents being attacked. Instead of targeting all the agents, we only attack a subset. This subset is not randomly chosen; rather, it is based on the trained network $f_{\theta}$’s loss. We select a subset $\mathcal{M}\subset\mathcal{N}$ of the agents, where $m=|\mathcal{M}|<n$ with the lowest loss J, as this indicates that most of the agents have aligned perceptions regarding the informations contained in the $m$ observations.
Formally the subset $\mathcal{M}$ is defined as follows:
$\mathcal{M}=\operatorname{argmin}_{\begin{subarray}{c}\mathcal{S}\subset\mathcal{N}\\ |\mathcal{S}|=m\end{subarray}}\sum_{i\in\mathcal{S}}\text{J}_{i}(\mathbf{o};\theta)$ (8)
where $\text{J}_{i}$ is the per-agent loss defined as:
$\text{J}_{i}(\mathbf{o};\theta)=\left(f_{\theta}(o_{-i})-o_{i}\right)^{2}$ (9)
IV-E Free misalignment attacks
Our goal is to leverage the alignment intuition without access to any internal information about the agents (e.g., observations, actions, policies, etc.). To illustrate our method, imagine a 2D game in which two agents must coordinate to catch an object in order to receive a reward. Each agent receives the $(x,y)$ coordinates of the goal object as its observation. To introduce misalignment into the agents’ observations, an adversary can attempt to push the agents in directions that are not aligned (not similar and different). In our simple example, an effective adversarial strategy is to push each agent in perpendicular directions (e.g., $\theta=90^{\circ}$ in Equation (10), similar to cosine similarity in natural language processing).
$\delta_{1}^{\top}\delta_{2}=\lvert\delta_{1}\rvert\lvert\delta_{2}\rvert\cos\theta=0$ (10)
This idea generalizes to high-dimensional observation spaces using vector orthogonality. Suppose we have $n$ agents, each receiving an observation of dimension $d$. To design a misalignment attack, we will generate a matrix $\boldsymbol{\delta}$:
$\boldsymbol{\delta}=\begin{bmatrix}\delta_{11}&\delta_{12}&\cdots&\delta_{1d}\\ \delta_{21}&\delta_{22}&\cdots&\delta_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ \delta_{n1}&\delta_{n2}&\cdots&\delta_{nd}\end{bmatrix}\in\mathbb{R}^{n\times d}$ (11)
That must satisfy two conditions:
  1. 1.
    Condition 1: The rows are orthogonal :
    $\delta_{i}^{\top}\delta_{j}=0\quad\forall i,j\in\mathcal{N},\ i\neq j,$ (12)
  2. 2.
    Condition 2: Satisfy the attack budget constraint
    $\|\delta_{i}\|_{\infty}\leq\epsilon\quad\forall i\in\mathcal{N}$ (13)
Generating such perturbations is non-trivial. A simple greedy approach that constructs $\boldsymbol{\delta}_{i}$ one at a time can easily satisfy the budget constraint (13), but it becomes difficult to guarantee orthogonality (12), especially in high-dimensional spaces. Another strategy is to sample a random matrix (e.g., from a normal distribution) and orthogonalize it using QR decomposition. While this guarantees orthogonality, it does not ensure the budget constraint is satisfied.
To generate perturbations that satisfy both constraints, we use partial Hadamard matrices. A full Hadamard matrix $\mathbf{H}\in\mathbb{R}^{d\times d}$ is a square matrix with entries in $\{+1,-1\}$ satisfying:
$\mathbf{H}\mathbf{H}^{\top}=d\,\mathbf{I}.$
By construction, its rows are orthogonal. A partial Hadamard matrix is obtained by selecting a subset of rows from a full Hadamard matrix. The resulting matrix $\tilde{\mathbf{H}}\in\mathbb{R}^{n\times d}$, with $n<d$, retains exact orthogonality between its rows.
It is clear that the following matrix satisfies both conditions (1) and (2):
$\boldsymbol{\delta}=\epsilon\times\tilde{\mathbf{H}}$ (14)
In our implementation, we generate Hadamard matrices using Sylvester’s construction. It is important to note that Hadamard matrices exist only for specific dimensions: when $d$ is a multiple of 4.
To bypass the requirement that $d$ must be a multiple of 4, we generate a full Hadamard matrix of size $\tilde{d}$, where $\tilde{d}$ is the largest power of two such that $\tilde{d}\leq d$:
$\tilde{d}=2^{\left\lfloor\log_{2}d\right\rfloor}$ (15)
We then pad the remaining $d-\tilde{d}$ columns with zeros. This padding affects neither orthogonality nor the budget constraint.
As an example when $n=3$ and $d=10$, we generate the following matrix in Figure 3: the first $\tilde{d}=8$ columns correspond to the scaled partial Hadamard matrix, while the remaining $d-\tilde{d}=2$ columns are zero-padded to extend the dimensionality to $d=10$.
Figure 3: Illustration: Padded partial Hadamard matrix for $n=3,d=10$.
We can theoretically combine the Hadamard attack with the Align attack in two ways: (1) by using a scaled partial Hadamard matrix to initialize the adversarial perturbations in Algorithm 2, line 2; or (2) by encouraging the perturbations of the Align attack to be orthogonal by adding a penalty to Equation (7):
$\lambda\left\|\boldsymbol{\delta}^{\top}\boldsymbol{\delta}-\boldsymbol{I}\right\|_{2}^{2}.$
We implemented these two approaches, but did not notice any substantial improvements. However, using the lightweight version of the Align attack from the previous section to select a subset of agents and injecting Hadamard perturbations proved very effective in our experiments.
V Experiments
TABLE II: QMIX Hyperparameters Used for SMAC and LBF benchmark
Hyperparameter SMAC LBF
Runner type parallel episode
Batch size per run 8 1
Network hidden dimension 128 64
Learning rate 0.001 0.0003
Standardise rewards True True
Use RNN True True
Epsilon anneal time 50000 200000
Target update interval 200 0.01
Training batch size 64 32
Maximum timesteps ($t_{max}$) 20,050,000 5050000
TABLE III: MAPPO Hyperparameters for SMAC benchmark.
Environments hdim1 lr Stand2 ent q_nstep Buffer Epochs
MMM2 64 0.0005 False 0.001 10 10 4
27m_vs_30m 64 0.0005 False 0.001 10 10 4
10m_vs_11m 128 0.001 True 0 5 8 10
SMACv2 128 0.0003 True 0.001 5 10 4
protoss_5_vs_5 64 0.0005 False 0.001 10 10 4
terran_5_vs_5 128 0.0003 True 0.001 5 10 4
zerg_5_vs_5 128 0.0003 True 0.001 5 10 4
1hdim: Network hidden dimension 2Stand: Standardise rewards
TABLE IV: MAPPO Hyperparameters for LBF benchmark.
Environment lr hdim Stand RNN q_nstep
8x8-3p-2f 0.0003 128 False False 5
8x8-3p-2f-coop 0.0005 32 True True 5
1s-8x8-3p-2f 0.0003 128 False True 10
2s-8x8-3p-2f 0.0003 128 False False 5
2s-8x8-3p-2f-coop 0.0005 128 True True 5
10x10-4p-2f 0.0003 128 False False 5
10x10-4p-2f-coop 0.0003 32 True True 5
1s-10x10-4p-2f 0.0003 128 False True 10
2s-10x10-4p-2f 0.0003 128 False False 5
2s-10x10-4p-2f-coop 0.0005 128 True False 5
TABLE V: MAPPO Hyperparameters for RWARE.
Environment lr Stand RNN q_nstep $t_{max}$
tiny-4ag 0.0005 False False 10 25M
tiny-4ag-sr2 0.0008 False False 10 25M
tiny-4ag-sr3 0.0008 False False 10 25M
small-4ag 0.0005 False False 10 25M
small-4ag-sr2 0.0008 False True 10 25M
small-4ag-sr3 0.0008 False False 20 40M
The goal of this section is to validate the effectiveness of our attack and identify the factors contributing to its outcome. To this end, we address the following key questions:
Question 1: Are the attacks effective? We seek to confirm that the execution of our attacks successfully compromises the performance of collaborative MARL.
Question 2: How do our attacks compare to standard baselines? We want to compare to standard adversarial attacks and see if our attacks outperform existing ones.
Question 3: Do the attacks generalize across different MARL algorithms? We examine whether the attacks affect both value-based and policy-based algorithms.
Question 4: What factors influence the Align attack’s outcome? We investigate several factors that could impact the Align attack’s success, primarily focusing on the size of the collected data, the architecture of the neural network $f_{\theta}$, the number of attacked agents, and the number of iterations $K$ of the PGD algorithm.
In the next section, we outline the experimental tools that help us to answer the aforementioned questions.
V-A Experimental setup
In order to answer our questions, we carefully choose multiple environments from three different MARL benchmarks:
  • Level-Based Foraging (LBF) [20]: LBF is a grid-world environment where agents collect randomly scattered food. We select three tasks: fully observable, highly cooperative, and partially observable. In fully observable tasks, agents receive the attributes of all entities in the grid (position and level). In the highly cooperative setting ( with the “-c” flag), food can only be collected if all agents act on it simultaneously. These setups provide overlapping observations, making alignment easier to measure. Partially observable tasks are expected to be more challenging for our attacks. In these tasks, agents only see a limited area around themselves. We test two settings: a 2-square radius (“-2s”) and a 1-square radius (“-1s”). We use 10 LBF environments in total. Details are in Table VI.
  • Multi-Robot Warehouse (RWARE) [20]: RWARE is a partially observable multi-agent environment where robots are tasked with collecting requested shelves inside a warehouse. We test with three different levels of partial observability, where agents observe a $7\times 7$, $5\times 5$, or $3\times 3$ grid. This benchmark also allows us to experiment with observations of higher dimensionality, in contrast to LBF’s observations. See table VII for details.
  • StarCraft Multi-Agent Challenge (SMAC) [3]: SMAC is the most widely used benchmark in c-MARL. It provides partially observable environments and includes scenarios with many agents, which helps us evaluate how our algorithms scale with a large number of agents. Since SMAC tasks are combat scenarios, they allow us to test our algorithms in situations where agents may become unavailable (e.g., due to death in SMAC). This mirrors many realistic settings where we might lose access to the target agents. See table VIII for details
TABLE VI: Selected LBF environments: we select (FO) fully observable, (PO) partial observable, and (HC) highly cooperative tasks.
Environment FO PO HC $\text{dim}(o_{i})$ N
8x8-3p-2f 15 3
8x8-3p-2f-coop 15 3
1s-8x8-3p-2f 15 3
2s-8x8-3p-2f 15 3
2s-8x8-3p-2f-coop 15 3
10x10-4p-2f 18 4
10x10-4p-2f-coop 18 4
1s-10x10-4p-2f 18 4
2s-10x10-4p-2f 18 4
2s-10x10-4p-2f-coop 18 4
TABLE VII: Selected RWARE environments: we select sensory range (SR) of $3\times 3$, $5\times 5$ and $7\times 7$.
Environment PO SR $\text{dim}(o_{i})$ N
tiny-4ag $3\times 3$ 71 4
tiny-4ag-sr2 $5\times 5$ 183 4
tiny-4ag-sr3 $7\times 7$ 351 4
small-4ag $3\times 3$ 71 4
small-4ag-sr2 $5\times 5$ 183 4
small-4ag-sr3 $7\times 7$ 351 4
TABLE VIII: Selected SMAC environments: all environments are partially observable (PO)
Environment PO $\text{dim}(o_{i})$ N
27m_vs_30m 285 27
10m_vs_11m 105 10
MMM2 176 10
protoss_5_vs_5 92 5
terran_5_vs_5 82 5
zerg_5_vs_5 82 5
As our paper considers test-time attacks, we need to first have well-trained agents. We train the agents in each environment with QMIX  [26] and MAPPO [29] using the hyperparameters specified in tables II, III, IV , and LABEL:tab:mappo_rware_hyper. We use the Epymarl library [20] for training. We only report hyperparameters that are different from the default values found in Epymarl configurations. For the RWARE benchmark, we only train MAPPO agents, as QMIX performs poorly at this task [20]. Once the training is complete, we freeze the policy weights.
For each environment, the adversary starts by collecting a dataset of observations from 5,000 environment transitions (i.e., $\mathcal{T}^{c}=5000$). This dataset is subsequently used to train the neural network $f_{\theta}$. We test with three different architectures:
  • Feedforward neural network: It consists of three hidden layers with 1024 hidden units.
  • Recurrent neural network: For LBF and SMAC, we use one recurrent layer with a hidden dimension of 1024. For RWARE, we use 2 consecutive layers.
  • Encoder-only transformer: For RWARE and LBF, we use d_model=256 and dim_feedforward=2048 with 16 heads. For SMAC, we use d_model=128 and dim_feedforward=1024 with 8 heads.
We use a learning rate of $0.0001$ and a batch size of 64. The feedforward and transformer networks are trained for 100 epochs, while the RNN network is trained for 300 epochs.
Adversarial baselines: We use the following baselines:
  • White box attack: It is the most powerful attack; it assumes access to the agent policy. In our paper, the white-box perturbation corresponds to the perturbation that minimizes the probability (or the q-value for QMIX) of the optimal action:
    $\delta=\arg\min_{\delta}\pi(\mathbf{o}+\delta,a^{*})\quad\text{s.t}\quad\|\delta\|_{\infty}\leq\epsilon$ (16)
  • Random attacks: Most previous works used only uniform distributions $\delta_{i}\sim\mathcal{U}(-\epsilon,\epsilon)$ and normal distributions $\delta_{i}\sim\mathcal{N}(0,\epsilon^{2})$. In addition to these two distributions, we add a non-symmetric distribution and a temporally correlated noise: for the former, we use the exponential distribution $\mathrm{Exp}(\lambda)$. The value of $\lambda$ is chosen such that 99% of the sampled values satisfy the budget constraint:
    $\lambda=-\frac{\ln(0.01)}{\epsilon}$ (17)
    As correlated noise, we use the Ornstein–Uhlenbeck (OU) process:
    $\delta_{t+1}=\delta_{t}+\theta(\mu-\delta_{t})\Delta t+\sigma\sqrt{\Delta t}\,\mathcal{N}(0,1)$ (18)
    In our experiments, we use:
    $\mu=\epsilon,\theta\approx\epsilon,\sigma\approx\frac{\epsilon}{10}\pm 0.001,\Delta t=1$
When presenting numerical results, and in order to avoid overloading the figures with curves of random attacks, we report for each $\epsilon$ the performance of the best random attack (i.e, the random noise with the lowest return), and it will be referred to as ”Random attack”.
Although it is not entirely fair to compare any adversarial attack with limited access to a white-box attack, we include the latter in our experiments to form an idea of the performance of a perfect-knowledge attack. We also use it to select appropriate values for the attack budget $\epsilon$, since it would be pointless to choose budgets where the random attacks outperform the white‑box attack.
We test different values of attack budget $\epsilon\in\{0.03,0.05,0.07,0.09,0.1,0.15,0.2\}$ for RWARE and SMAC and $\epsilon\in\{0.1,0.15,0.2,0.25,0.5,0.75\}$ for LBF. Throughout all experiments, we maintain a fixed number of PGD iterations at $K=10$ and the step size of $\alpha=\frac{\epsilon}{K}$.
Metrics: We use the episodic return to evaluate our attacks. In order to report robust results, we report the Interquartile Mean (IQM) [1] of the returns from 50 independent episodes, and we also report the 95% confidence intervals (CIs). IQM is robust to outliers and has lower bias than the median [1]. Our code is open-source.  [^1] Github link: https://github.com/AmineAndam04/black-box-marl
V-B Numerical results
Figure 4: Performance on SMAC: IQM returns and 95% CIs estimated using 50 runs.
Figure 5: Performance on RWARE: IQM returns and 95% CIs estimated using 50 runs.
Figure 6: Performance on LBF: IQM returns and 95% CIs estimated using 50 runs.
In the following, we provide answers to our four questions. Unless otherwise stated, the presented figures and tables correspond to attacks executed against agents trained with MAPPO and use RNNs to train $f_{\theta}$ for the Align attack. We start by answering the first two questions.
Figures 45, and 6 present the IQM returns (y-axis) for SMAC, RWARE, and LBF benchmarks, respectively, across selected $\epsilon$ values (x-axis). The figures show the performance of our proposed algorithms ’Align’ and ’Hadamard’, as well as the white-box and random noise baselines. The values at $\epsilon=0$ represent the benign performance of the agents during deployment.
- LBF: In fully observable scenarios (the first two rows in Figure 6), our attacks significantly impact the agents’ performance, particularly in highly cooperative tasks where observations overlap. In such scenarios, the returns drop to very low values more rapidly. The increasing level of cooperation also has the same effects in partially observable tasks: comparing the third and fourth rows, which represent the same task with the latter being highly cooperative, the returns drop to much lower values at lower attack budgets. Surprisingly, it’s noticeable that the Hadamard attack matches or outperforms the Align attack in highly cooperative and partially observable scenarios (the last two columns). We notice the same pattern with the RWARE benchmark.
- RWARE: In general, both of our attacks perform well across the tasks. In scenarios with the highest degree of partial observability (sr $3\times 3$, columns 1 and 4 in Figure  5), while the align attack still performs well, it’s slightly outperformed by the Hadamard attack. As the partial observability decreases, the alignment network $f_{\theta}$ becomes more useful. The results on the RWARE benchmark also show us that our attacks are not negatively impacted by high-dimensional observations (see Table VII).
- SMAC: The Align attack maintains a consistent performance across the SMAC games. The Hadamard attack, on the other hand, struggles to perform well or significantly outperforms random noise attacks in three games: 10m_vs_11m, 27m_vs_30m, and protoss_5_vs_5. Besides the partial observability and high dimensionality, the SMAC benchmark allows us to confirm that the Align attack performs well on systems with a high number of agents (the first three columns) and in situations with no persistent access to the agents (all SMAC games).
Regarding the performance of the lighter version of the Align attack, we present the IQM return curves in Figures 78, and 9. In each figure, we plot the return curves with a decreasing number of targeted agents. For the align attack, we use $f_{\theta}$ to select the most vulnerable agents, as explained in Section IV-D. For the Hadamard attack, agents are selected randomly. Based on the results, we make two observations: First, the Align attack generally maintains a good performance in scenarios where 50% or more of the agents are targeted. Second, the Align attack is less affected by the reduction in the number of targeted agents than the Hadamard attack. To illustrate this, we consider tasks where Hadamard outperformed Align attack: In RWARE we take “small-4ag” and “tiny-4ag” (1st and 4th columns in Figure 8); in LBF we take the four tasks in the last two columns of figure 9: “2s-8x8-coop”, “2s-10x10-coop”, “1s-8x8”, and “1s-10x10”. In all these six tasks, Hadamard outperforms the Align attack when all agents are under attack, but as fewer agents are targeted, we notice the opposite trend: the Align attack performs better.
The last observation also demonstrates that the Align attack effectively selects the most vulnerable agents to target, whereas the Hadamard attacks lack such a mechanism. To further test the potential of using the idea behind the Align attack as a target selection method, we combine both attacks: we first select the most vulnerable agents using the Align network (as in Equation 8), and then, instead of injecting the Align perturbation, we inject Hadamard perturbations.
In Tables IXX, and  XI we report the additional percentage drop in IQM returns when adding the selection mechanisms to the Hadamard attack ($m$ refers to the number of targeted agents): for example, if the Hadamard attack alone results in a -15% return drop relative to the benign return, and the reported value in one of the tables is -10%, this means that the return drop for the targeted Hadamard attack is -25% (-15 + (-10) %) relative to the benign return. We report results of the six tasks mentioned above, as well as of partially observable tasks. Across all benchmarks, the targeted Hadamard attack significantly outperforms the untargeted version, achieving a maximum additional drop of -57%. On average, the additional drop is -11.5% on the LBF benchmark, -6.18% on RWARE tasks, and -6.28% on SMAC games. These results confirm the effectiveness of the Align attack for target selection.
Figure 7: Performance of lighter attacks on SMAC: IQM returns and 95% CIs curves estimated using 50 episodes. For games with 10 agents, we report return curves when attacking 10, 7,5,3, and 1 agent(s). For games with 5 agents, we show the returns when targeting 5,3,2,1 agent(s). For the 27m_vs_30m map, we report when having 27,20 15,10, and 5 victims.
Figure 8: Performance of lighter attacks on RWARE: IQM returns and 95% CIs curves estimated using 50 episodes. We report return curves when attacking 4,3,2,3, and 1 agent(s).
Figure 9: Performance of lighter attacks on LBF: IQM returns and 95% CIs curves estimated using 50 episodes. For games with 3 agents, the first row, we report return curves when attacking 3, 2, and 1 agent(s). For tasks with 4 agents, the second row, we include the returns when targeting 4,3,2,1 agent(s).
TABLE IX: Targeted Hadamard attack on LBF: additional drop in IQM returns of the targeted attack relative to untargeted Hadamard attack
Task $m$ $\epsilon=0.15$ $\epsilon=0.2$ $\epsilon=0.25$
1s-10x10-4p-2f 2 -23.57 -4.99 -6.6
3 2.32 -6.28 -7.48
2s-10x10-4p-2f 2 -10.02 -21 -8,57
3 -10.49 -11.71 -15.26
2s-10x10-4p-2f-coop 2 -3.84 -15.36 -17.30
3 -3.84 -13.46 -15.38
1s-8x8-3p-2f 1 -6.01 -0.96 -19.24
2 -7.42 -0.79 -19.34
2s-8x8-3p-2f 1 -0.93 3.15 1.88
2 3.75 -18.10 -11.17
2s-8x8-3p-2f-coop 1 0 -3.84 -3.84
2 -32.69 -48.07 -57.69
TABLE X: Targeted Hadamard attack on RWARE: additional drop in IQM returns of the targeted attack relative to untargeted Hadamard attack
Task $m$ $\epsilon=0.05$ $\epsilon=0.09$ $\epsilon=0.15$
tiny-4ag 1 0.22 -0.37 -4.50
2 0.52 -1.12 -9.97
3 -1.35 -2.1 -16.80
small-4ag 1 -9.80 -6.71 -2.35
2 3.26 -11.61 -12.70
3 -2.17 -14.51 -19.23
TABLE XI: Targeted Hadamard attack on SMAC: additional drop in IQM returns of the targeted attack relative to untargeted Hadamard attack
Task $m$ $\epsilon=0.05$ $\epsilon=0.09$ $\epsilon=0.15$
10m_vs_11m 3 -8.3 -10.17 -3.71
7 -14.35 0.40 1.98
MMM2 3 3.26 0.73 -6.32
7 -2.76 1.22 -7.10
protoss_5_vs_5 2 -4.40 -7.58 -4.10
3 -10.10 -1.37 -4.88
terran_5_vs_5 2 4.05 -3.05 -19.48
3 -12.35 -21.03 -5.15
zerg_5_vs_5 2 -22.28 -11.09 -3.77
3 0.27 -12.86 -4.40
V-C Ablations
Trained agents: In the following paragraphs, we address the remaining two questions. In Figures 10 and 11, we compare the performance of the Align attack against MAPPO and QMIX agents. To normalize the results across tasks, the y-axis shows the percentage drop of the return caused by injecting adversarial perturbations, relative to the benign rewards. Both figures show that our approach works for both policy-based and value-based algorithms. On the LBF benchmark, the performance is nearly identical. For SMAC, however, the Align attack tends to be more impactful on QMIX agents at higher $\epsilon$ values.
Figure 10: Performance across MARL algorithms on LBF: IQM returns and 95% CIs estimated across 50 runs and averaged across all LBF tasks.
Figure 11: Performance across MARL algorithms on SMAC: IQM returns and 95% CIs estimated across 50 runs and averaged across all SMAC tasks.
Figure 12: Collected data size effect across MARL benchmarks: IQM return drop relative to benign performance, averaged across tasks for each benchmark. Data size values are reported in thousands ($\times 1000$)
Data collection: In Figure12, we evaluate the Align attack while varying the number of collected data used to train the neural network $f_{\theta}$. We examined datasets ranging from collecting 1,000 steps (1k, in blue) to 100,000 steps (100k, in brown). Surprisingly, the results show that only a few samples are needed for a successful attack, as there is no noticeable difference when using larger datasets. This provides a significant advantage to the adversary: collecting large amounts of data may compromise the stealthiness of the attack and can be costly. Furthermore, requiring less data reduces the storage and computational burden during training, significantly lowering the costs of executing the attack.
Network architecture: Figure 13 compares the performance of the Align attack across different neural architectures. Across all three benchmarks, the RNN architecture often performs best, especially on the RWARE benchmark, where the gap between the RNN and the other networks is significant. The performance of the MLP and the Transformer network is comparable. However, given that MLPs are faster to train than the Transformer (as the MLP only has 3 hidden layers), they offer a better cost-performance trade-off. RNNs also have this advantage, as we use shallow variants with only one or two recurrent layers. Although we experimented with deeper and more complex neural architectures, we did not notice any noticeable improvements in performance.
Figure 13: Neural architecture effect across MARL benchmarks: IQM return drop relative to benign performance, averaged across tasks for each benchmark.
PGD and FGSM: We investigate the impact of the number of steps $K$ used in our algorithm. We are particularly interested in the scenario where $K=1$, which corresponds to using FGSM, to assess whether PGD is necessary for effective performance. Figure 14 compares the performance for different values of $K\in\{1,5,10\}$. A notable observation is that FGSM alone is sufficient to achieve satisfactory results. This is very important from a practical perspective: an increasing number of iterations means that we need more time to generate the perturbations, whereas a single-step method like FGSM enables fast computation. Given that 1 iteration is enough, this will help us implement faster attacks, especially in real-time scenarios.
Figure 14: PGD’s iteration effect across MARL benchmarks: IQM return drop relative to benign performance, averaged across tasks for each benchmark.
V-D Discussion
In most of the adversarial MARL literature, and as in our experimental section, the effectiveness of an adversarial attack is primarily measured using the episodic return. While this is a valid and essential metric, it should not be the sole focus. Other metrics, especially in real-world scenarios, are equally important. One such metric that we believe is often overlooked despite its significance is the episode length, particularly when the adversary has constrained capabilities, as is the case in our scenario.
In these settings, it is unrealistic to expect that a weak adversary will fully incapacitate the multi-agent system. A more plausible outcome would be delaying task completion, i.e., increasing the episode length, even if the tasks are eventually completed. This is a critical consideration, as prolonged task completion can result in significant resource overhead by increasing the consumption of resources like electricity and computation. Therefore, from the adversary’s perspective, increasing operational costs may be considered a satisfactory outcome.
To illustrate this point, we examine the Align attack on the LBF benchmark. In table XII, we report the percentage increase in episode lengths relative to the normal episode length. We notice that the Align attack significantly affects episode lengths, with increases reaching at least two digits in most tasks, and up to 285% in the worst case.
TABLE XII: Episode length increase (%) on LBF tasks
Tasks 0.1 0.15 0.2 0.25 0.5 0.75
1s-8x8 1.30 1.55 2.09 3.63 12.09 12.18
1s-10x10 -5.68 -9.66 2.26 7.94 16.64 22.88
2s-8x8-coop 9.59 43.55 138.52 152.49 170.86 170.86
2s-8x8 5.17 18.84 31.93 44.26 72.24 72.24
2s-10x10-coop 34.44 23.78 65.78 68.52 79.04 85.19
2s-10x10 12.45 -2.72 14.36 20.78 52.38 54.29
8x8-coop 10.59 29.67 103.86 188.56 226.80 226.80
8x8 5.94 21.06 51.93 75.93 193.21 285.80
10x10-coop 23.49 18.89 23.43 42.13 190.87 214.86
10x10 0.66 20.32 11.93 29.41 49.43 110.93
Limitations: Our proposed algorithm faces the same core challenges as any MARL training procedure, such as partial observability, which we discuss thoroughly in the paper. In addition, issues like handling heterogeneous agents and multi-modal observations may further affect the effectiveness of our attack. We aim to address these limitations in future work.
VI Conclusion
In this paper, we present two novel adversarial attacks specifically designed for collaborative MARL during deployment. We consider a black-box scenario under highly constrained conditions: First, a scenario where the adversary has access solely to the observations of deployed agents. Second, a more constrained scenario where the adversary has no access at all. Our approach exploits the reliance of collaborative agents on aligned perceptions for effective cooperation and adds small perturbations to the agents’ observations specifically designed to induce misaligned perceptions.
Our investigation of key factors impacting the Align attack provides us with valuable insights into the computational efficiency of our algorithm. Notably, our method requires minimal data to generate effective perturbations; we found that as few as 1,000 steps are sufficient to create impactful damage. Moreover, the adversary can choose to execute the attack in fewer steps while still achieving satisfactory results.
Our work answers the question of whether and how an adversary with very limited information can significantly undermine the performance of collaborative MARL systems. At first glance, it may seem unlikely that an adversary with such limited access could inflict any significant damage. Yet, as our findings demonstrate, an adversary with minimal access can severely undermine these collaborative systems. Just as important is the question of why c-MARL systems are vulnerable to such attacks. Understanding why this vulnerability exists is critical for designing defensive strategies and robust MARL algorithms. A natural direction for future work is to develop robust defense mechanisms, either by detecting such attacks or designing algorithms that are inherently resilient.
Finally, our goal is to draw the attention of the research community to previously unknown vulnerabilities present in existing collaborative c-MARL algorithms. These unexpected vulnerabilities must question our assumptions regarding the security and reliability of multi-agent systems. Moreover, such work will enhance the transparency about the security implications of intelligent systems, which is essential for the widespread adoption of such systems.